A positron of mass m moving with speed Vo enters into a region of uniform magnetic field B ,(produce out of the plane of paper) . The width of the region of magnetic field is " (3mVo)/(2eB)". Find the time spent by the particle in magnetic field if "theta" is nearly 90°.
(a) (2πm)/(3Be). (b). (3πm)/2Be
(c) (2πm)/eB. (d). Infinite
1
Expert's answer
2019-03-19T12:21:01-0400

The law of motion is mv2/r = eBv, hence mω = eB, that is, the angle θ = π/2 = ωτ, therefore τ = π/2ω = πm /2eB.